# Spectuner: A Framework for Automated Line Identification of Interstellar Molecules

### v1

@@ Abstract
Interstellar molecules, which play an important role in astrochemistry, are identified using observed spectral lines. Despite the advent of spectral analysis tools in the past decade, the identification of spectral lines remains a tedious task that requires extensive manual intervention, preventing us from fully exploiting the vast amounts of data generated by large facilities such as ALMA. This study aims to address the aforementioned issue by developing a framework of automated line identification. We introduce a robust spectral fitting technique applicable for spectral line identification with minimal human supervision. Our method is assessed using published data from five line surveys of hot cores, including W51, Orion-KL, Sgr B2(M), and Sgr B2(N). By comparing the identified lines, our algorithm achieves a recall of $\sim 84\%-98\%$. Our code, named spectuner, is publicly available on GitHub.
@@ 1 Introduction
The main goal of this study is to develop an automated line identification framework of interstellar species. Our algorithm, which builds on xclass (Möller et al., 2017), can automatically analyze an observed single-pointing spectrum and produce the best-fitting spectrum, associated parameters, and a table of identified lines with minor human intervention. Our algorithm is assessed using published line survey data of hot cores towards W51, Orion KL, Sgr B2(M) and Sgr B2(N). Our tool, named spectuner, is publicly available on GitHub [^1] https://github.com/yqiuu/spectuner.
@@ 2.1 Overview
We introduce the spectral line model and the spectroscopic database in Section 2.2. Our fitting techniques are detailed in Sections 2.3 and 2.4. The greedy method and the criteria for combining a fitting results are described in Section 2.5.
@@ 2.2 The spectral line model
This study employs xclass (Möller et al., 2017) to generate model spectra. The xclass code predicts model spectra by solving the one-dimensional radiative transfer equation for an isothermal object, assuming local thermodynamic equilibrium (LTE). In this study, we assume that the spectra of different molecules combine linearly, without considering any absorption features. The spectroscopic data are sourced from the Virtual Atomic and Molecular Data Center (VAMDC) (Endres et al., 2016), which contains entries from the Cologne Database for Molecular Spectroscopy (CDMS) (Müller et al., 2001, 2005) and the Jet Propulsion Laboratory (JPL) database (Pickett et al., 1998).
In this work, we fit the spectrum of a molecule, encompassing all possible states within the specified frequency range. Our code is capable of automatically selecting these possible states from the database.
In XCLASS, assuming a single component, the spectrum of one molecule of one state is characterized using five parameters: source size $\theta$, excitation temperature $T_{\text{ex}}$, column density $N_{\text{tot}}$, velocity width $\Delta v$, and velocity offset $v_{\text{LSR}}$. We assume that source size, velocity width and velocity offset are shared among different states, while the other parameters are independent. This approach reduces the degrees of freedom, thereby simplifying the problem. All molecules considered in this work are assumed to consist of only one component for simplicity. In addition, we do not account for any absorption features in the spectrum. Frequency channels that are below the background temperature are set to this temperature.
@@ 2.4 The peak matching loss function
Figure 3: Visualization of $S_{\text{TP}}^{i}$ and $S_{\text{FP}}^{i}$ defined in Equations 18 to 22. Black solid lines illustrate the observed spectrum of W51-Mopra (see Section 3.1). Red lines show the best-fitting spectra of 13CH3OH, C2H5CN and CH3OCH3 obtained with the peak matching loss in the combined phase. The complete identification results for W51-Mopra are presented in Section 4. Each black (blue) vertical line represents a true (false) positive peak identified using our peak finding algorithm as described in Section 2.4.
@@
The first term is known as the mean absolute error (MAE), which is less sensitive to outliers and is commonly used in machine learning context. It is defined as
@@
where $N$ is the number of observed frequency channels.
@@
$\mathcal{L}_{\text{TP}}=\begin{cases}D_{\text{TP}}-f_{\text{dice}}I^{\text{obs}},&\,I^{\text{obs}}\geq I^{\text{pred}},\\ \min(D_{\text{TP}}-f_{\text{dice}}I^{\text{obs}},0),&\,I^{\text{obs}}<I^{\text{pred}},\\ \end{cases}$ (5)
@@
where $d_{\cap}$, $d_{\text{obs}}$, and $d_{\text{pred}}$ are shown in the lower panel of Figure 2. The use of $f_{\text{dice}}$ has two effects. First, it smooths the transition between the cases with and without intersection. Secondly, it could measure the similarity in shape between two peaks. When computing the integrals in Equations 7, 8, and 9, the trapezoidal rule is adopted. Within the interval, we only use channels given by the observed spectrum, without including any additional points, while linear interpolation is applied at the boundaries. Equation 2 can yield negative values when true positive peaks (TPPs) are identified, signifying a strong correspondence between the peaks in the observed and model spectra.
@@
where $P$ is the value of the prominence and $\nu_{\text{FP}}$ is the central frequency of the false positive peak. $\nu_{\text{left}}$ and $\nu_{\text{right}}$ are the central frequencies of the nearest observed peaks on the left and right sides respectively. The trapezoidal rule is adopted to compute Equations 12 and 13. If a false positive peak is hidden in a broad observed peak, e.g. the left false positive peak in the lower panel of Figure 2, there is no need to penalize it, and therefore we introduce a maximum operation in Equations 12 and 13. The use of $D_{2}$ is to smooth the transition between the cases if a local maximum is identified and not identified as a peak. $W(\nu)$ allows to reduce $\mathcal{L}_{\text{FP}}$ by shifting the central frequency in addition to decreasing $T^{\text{pred}}_{\nu}$.
@@ 2.5 Combining individual fitting results
This work proposes a greedy algorithm to combine individual fitting results (see Algorithm 1). In the algorithm, a ”pack” consists of a molecule list, model parameters and model spectrum, which can initiate a new fitting. The algorithm input is a list of packs, each constructed from individual fitting results. The algorithm outputs a combined result and fitting results of some candidate molecules which do not satisfy the criteria described in the following paragraph.
@@
where $I_{\text{pred}}^{i}$, $D_{\text{TP}}^{i}$, and $D_{\text{FP}}^{i}$ are defined in Equations 9, 7, and 11. These quantities here are computed for each molecule. We note that a true (or false) positive peak can be contributed by multiple molecules, and we use $F_{\text{TP}}^{i}$ and $F_{\text{FP}}^{i}$ to take into account the different contributions. In the calculation of $F_{\text{TP}}^{i}$ and $F_{\text{FP}}^{i}$, contributions less than $5\%$ are excluded beforehand, resulting in renormalized fractions. Figure 3 provides a visualization of $S_{\text{TP}}^{i}$ and $S_{\text{FP}}^{i}$.
@@
We point out that a more straightforward method for combining all individual fitting results could involve simultaneously fitting the spectrum of all molecules. However, this method is considered to be impractical. The large number of molecules involved in the fitting task results in an excessive number of free parameters ($\sim$ 100 to 500), which requires considerable iterations for the optimizer to attain convergence. In contrast, our greedy algorithm is more efficient. Furthermore, due to the intricacy of the spectral lines, there are various choices for combining the individual fitting results, and our algorithm offers an interpretable solution.
@@ 3.1 W51-Mopra
We utilize the published data from Watanabe et al. (2017), who conducted a line survey towards the W51 region using the Mopra 22 m telescope. The survey covers the frequency ranges from 85.2 - 101.1 GHz and 107.0 - 114.9 GHz. The spectral data originate from the hot cores e1/e2, located at galactic coordinate ($l$, $b$) = (49.4898, -0.3874). Watanabe et al. (2017) identified 234 emission lines. The faintest line has a peak temperature of 60 mK, detected at $\sim 4.5\sigma$ level.
@@ 3.2 OrionKL-GBT
This study relies on the data published by Frayer et al. (2015), who conducted a line survey towards Orion KL. The observations were carried out using the Green Bank Telescope (GBT) and were centered on the celestial coordinates R.A. (J2000) 05:35:14, decl. (J2000) 05:22:27.5. The frequency range covered by the survey spans from 67.0 GHz to 93.6 GHz. Frayer et al. (2015) only identified lines with intensities brighter than 1 K, and in total, 140 emission lines were identified.
@@ 3.3 OrionKL-Tianma
We incorporate the data from Liu et al. (2022), who carried out a line survey targeting the Orion KL region using the Tianma 65 m radio telescope. The spectral data are centered on the celestial coordinates R.A. (J2000) 05:35:14.55, decl. (J2000) -05:22:31.0. The frequency coverage spans from 34.5 GHz to 50.0 GHz. A total of 597 emission lines were identified, including 177 classified as radio recombination lines. The faintest line detected has a peak intensity of 2.8 mK.
@@ 3.4 SgrB2M-IRAM and SgrB2N-IRAM
The most complex test data utilized in this study originate from Belloche et al. (2013), who presented a line identification analysis for both Sgr B2(M) and B2(N) within the frequency range of 80.0 GHz to 116.0 GHz. The observations were conducted using the IRAM 30 m telescope, targeting the positions at R.A. (J2000) 17:47:20s.4, decl. (J2000) -28:23:07.0 for Sgr B2(M), and R.A. (J2000) 17:47:20.0 and decl. (J2000) -28:22:19.0 for and B2(N), respectively. These sources are line rich. Belloche et al. (2013) identified 945 and 3675 emission lines in Sgr B2(M) and B2(N) respectively.
@@
Name Frequency coverage Line density Recall ($\chi^{2}$) Recall (PM) Duration (PM)
(GHz) (GHz-1) (hour)
W51-Mopra 85.2 - 101.1, 107.0 - 114.9 13.8 78% (118/152) 91% (139/152) 8.6
OrionKL-GBT 67.0 - 93.6 20.3 86% (90/105) 98% (103/105) 5.6
OrionKL-Tianma 34.5 - 50.0 30.5 59% (140/237) 87% (207/237) 14.8
SgrB2M-IRAM 80.0 - 116.0 38.0 60% (276/463) 85% (392/463) 25.6
SgrB2N-IRAM 80.0 - 116.0 59.2 47% (498/1055) 84% (881/1055) 59.0
Figure 4: Comparison of the best-fitting spectra using different loss functions. Black solid line illustrates the observed spectrum of SgrB2M-IRAM (see Section 3.4), and blue dashed lines represent the spectral lines of C2H3CN identified by Belloche et al. (2013). Blue and red lines demonstrate the best-fitting results achieved with the $\chi^{2}$ and peak matching loss functions respectively. The peak matching loss function is defined in Section 2.4. We compute and display both the $\chi^{2}$ and peak matching loss values corresponding to each best-fitting spectrum. These values imply that using the $\chi^{2}$ loss is inappropriate for spectral fitting of a single molecule.

### v2

@@ Abstract
Interstellar molecules, which play an important role in astrochemistry, are identified using observed spectral lines. Despite the advent of spectral analysis tools in the past decade, the identification of spectral lines remains a tedious task that requires extensive manual intervention, preventing us from fully exploiting the vast amounts of data generated by large facilities such as ALMA. This study aims to address the aforementioned issue by developing a framework of automated line identification. We introduce a robust spectral fitting technique applicable for spectral line identification with minimal human supervision. Our method is assessed using published data from five line surveys of hot cores, including W51, Orion-KL, Sgr B2(M), and Sgr B2(N). By comparing the identified lines, our algorithm achieves an overall recall of $\sim 74\%-93\%$, and an average precision of $\sim 78\%-92\%$. Our code, named spectuner, is publicly available on GitHub.
@@ 1 Introduction
The main goal of this study is to develop an automated line identification framework of interstellar species. Our algorithm, which builds on xclass (Möller et al., 2017), can automatically analyze an observed single-pointing spectrum and produce the best-fitting spectrum, associated parameters, and a table of identified lines with minor human intervention. Our algorithm is assessed using published line survey data of hot cores towards W51, Orion KL, Sgr B2(M) and Sgr B2(N).
@@
Our code, named spectuner, is publicly available on GitHub [^1] https://github.com/yqiuu/spectuner under a BSD 3-Clause License and version 1.0.0-beta is archived in Zenodo (Qiu et al., 2024).
@@ 2.1 Overview
We introduce the spectral line model and the spectroscopic database in Section 2.2. Our fitting techniques are detailed in Sections 2.3 and 2.4. The greedy method and the criteria for combining a fitting results are described in Section 2.5. A summary of the configuration parameters used in our algorithm is presented in Table 2.
@@ 2.2 The spectral line model
This study employs xclass (Möller et al., 2017) to generate model spectra. The xclass code predicts model spectra by solving the one-dimensional radiative transfer equation for an isothermal object, assuming local thermodynamic equilibrium (LTE):
$T_{\nu}=I^{\text{bg}}_{\nu}-J^{\text{CMB}}_{\nu}\\ +\sum_{m}\sum_{c}\eta(\theta^{m,c})(J_{\nu}(T^{m,c}_{\text{ex}})-I^{\text{bg}}_{\nu})(1-e^{-\tau^{m,c}_{\nu}}),$ (1)
where the sums run over the indices $m$ for different species and their components $c$. Here, $I^{\text{bg}}_{\nu}$ is the background intensity, $J^{\text{CMB}}_{\nu}$ is the cosmic background intensity, $\eta(\theta^{m,c})$ is the beam filling factor, and $\theta^{m,c}$ is the source size, $T^{m,c}_{\text{ex}}$ is the excitation temperature. The source function $J_{\nu}(T)$ follows the black body radiation. Equation 1 assumes that the spectra of different species combine linearly.
The optical depth $\tau^{m,c}_{\nu}$ is defined by
$\tau^{m,c}_{\nu}=\sum_{t}\tau^{m,c,t}_{\nu},$ (2)
with
$\tau^{m,c,t}_{\nu}=\frac{c^{2}}{8\pi\nu^{2}}N^{m,c}_{\text{tot}}\phi(\nu;\nu^{t},v^{m,c}_{\text{LSR}},\Delta v^{m,c})\\ \times\frac{g^{t}_{\text{u}}A^{t}_{\text{ul}}}{Q^{m}(T^{m,c}_{\text{ex}})}e^{-E^{t}_{\text{l}}/T^{m,c}_{\text{ex}}}\left(1-e^{-h\nu^{t}/T^{m,c}_{\text{ex}}}\right),$ (3)
where $N_{\text{tot}}$ is the column density, $v_{\text{LSR}}$ is the velocity offset, $\Delta v$ is the velocity width, and $\phi(\nu)$ follows a Gaussian profile. The spectroscopic data used to compute the optical depth, i.e. $\nu^{t}$,$g^{t}_{\text{u}}$, $A^{t}_{\text{ul}}$, $E^{t}_{\text{l}}$, and $Q^{m}(T)$, are from the Virtual Atomic and Molecular Data Center (VAMDC) (Endres et al., 2016), which contains entries from the Cologne Database for Molecular Spectroscopy (CDMS) (Müller et al., 2001, 2005) and the Jet Propulsion Laboratory (JPL) database (Pickett et al., 1998). Readers are referred to Möller et al. (2017) for a more detailed description for computing the model spectra.
In this work, we fit the spectrum of a molecule, encompassing all possible states within the specified frequency range. Our code is capable of automatically selecting these possible states from the database. For a single component, the spectrum of one molecule of one state is characterized using five parameters: source size $\theta$, excitation temperature $T_{\text{ex}}$, column density $N_{\text{tot}}$, velocity width $\Delta v$, and velocity offset $v_{\text{LSR}}$. We assume that the source size, velocity width and velocity offset are shared among different states, while the other parameters are independent. This approach reduces the degrees of freedom, thereby simplifying the problem.
In addition, this study make the following assumptions:
  1. 1.
    All molecules and isotopologues considered in this work are assumed to consist of only one component for simplicity.
  2. 2.
    We do not account for any absorption features in the spectrum. Frequency channels below the background temperature are set to this temperature.
@@ 2.3 Optimization
Table 2: Summary of the configuration parameters of the proposed automated line identification algorithm.
Symbol Value Description Section
$N_{\text{swarm}}$ 28a Number of particles used in the particle swarm optimization. Section 2.3
$N_{\text{trail}}$ 3 Number of trails to run the optimization algorithm. Section 2.3
$R$ 0.25 Relative height to compute peak widths. Section 2.4
[…]
$S_{X}$ $S_{3}>0.8$ Threshold of the $X$-th largest score to combine a fitting result. Section 2.5
$S_{\text{tot}}$ $S_{\text{tot}}>2.7$ Threshold of the total score to combine a fitting result. Section 2.5
@@ 2.4 The peak matching loss function
Figure 3: Visualization of $S_{\text{TP}}^{i}$ and $S_{\text{FP}}^{i}$ defined in Equations 21 to 25. The black solid lines illustrate the observed spectrum of W51-Mopra (see Section 3.1). The red lines show the best-fitting spectra of 13CH3OH and C2H5CN obtained with the peak matching loss in the combined phase. The complete identification results for W51-Mopra are presented in Section 4. Each black (blue) vertical line represents a true (false) positive peak identified using our peak finding algorithm as described in Section 2.4.
@@
The first term is known as the mean absolute error (MAE), which is defined as
@@
where $N$ is the number of observed frequency channels. The MAE is commonly used in statistics and machine learning due to its robustness against outliers (Klebanov et al., 2009). In addition, it can serve as a regularization term in regression models (Tibshirani, 1996).
@@
$\mathcal{L}_{\text{TP}}=\begin{cases}D_{\text{TP}}-f_{\text{dice}}I^{\text{obs}},&\,I^{\text{pred}}\geq I^{\text{obs}},\\ \min(D_{\text{TP}}-f_{\text{dice}}I^{\text{obs}},0),&\,I^{\text{pred}}<I^{\text{obs}},\\ \end{cases}$ (8)
@@
where $d_{\cap}$, $d_{\text{obs}}$, and $d_{\text{pred}}$ are shown in the lower panel of Figure 2. The use of $f_{\text{dice}}$ has two effects. First, it smooths the transition between the cases with and without intersection. Secondly, it could measure the similarity in shape between two peaks. When computing the integrals in Equations 10, 11, and 12, the trapezoidal rule is adopted. Within the interval, we only use channels given by the observed spectrum, without including any additional points, while linear interpolation is applied at the boundaries. Unlike a typical loss function such as the MAE, $\mathcal{L}_{\text{TP}}$ can produce negative values when the peaks in the observed and model spectra match well.
@@
where $P$ is the value of the prominence and $\nu_{\text{FP}}$ is the central frequency of the false positive peak. $\nu_{\text{left}}$ and $\nu_{\text{right}}$ are the central frequencies of the nearest observed peaks on the left and right sides respectively. The trapezoidal rule is adopted to compute Equations 15 and 16. If a false positive peak is hidden in a broad observed peak, e.g. the left false positive peak in the lower panel of Figure 2, there is no need to penalize it, and therefore we introduce a maximum operation in Equations 15 and 16. The use of $D_{2}$ is to smooth the transition between the cases if a local maximum is identified and not identified as a peak. $W(\nu)$ allows to reduce $\mathcal{L}_{\text{FP}}$ by shifting the central frequency in addition to decreasing $T^{\text{pred}}_{\nu}$.
@@ 2.5 Combining individual fitting results
This work proposes a greedy algorithm to combine individual fitting results (see Algorithm 1). In the algorithm, a ”pack” consists of a molecule list, model parameters and a model spectrum, which can initiate a new fitting. The algorithm input is a list of packs, each constructed from individual fitting results. The basic idea of the algorithm is that fitting results with lower values can better match more features in the observed spectrum. The algorithm initially sorts the individual fitting results by their loss values, then combines them sequentially, beginning with the one having the lowest loss value. A new optimization will be performed if there is line blending. The algorithm outputs a combined result and fitting results of some candidate molecules which do not satisfy the criteria described in the following paragraph.
@@
where $I_{\text{pred}}^{i}$, $D_{\text{TP}}^{i}$, and $D_{\text{FP}}^{i}$ are defined in Equations 12, 10, and 14. These quantities here are computed for each molecule. We note that a true (or false) positive peak can be contributed by multiple molecules, and we use $F_{\text{TP}}^{i}$ and $F_{\text{FP}}^{i}$ to take into account the different contributions. In the calculation of $F_{\text{TP}}^{i}$ and $F_{\text{FP}}^{i}$, contributions less than $5\%$ are excluded beforehand, resulting in renormalized fractions. Figure 3 provides a visualization of $S_{\text{TP}}^{i}$ and $S_{\text{FP}}^{i}$.
@@
We point out that a more straightforward method for combining all individual fitting results could involve simultaneously fitting the spectrum of all molecules. However, this method is not adopted in this study. The large number of molecules involved in the fitting task results in an excessive number of free parameters ($\sim$ 100 to 500), posing a significant challenge. El-Abd et al. (2024) conducted a spectral line analysis for NGC 6334I, who simultaneously fit the spectrum of 21 molecules. They fixed the excitation temperature and line width of each molecule to reduce the degrees of freedom. Furthermore, due to the intricacy of the spectral lines, there are various choices for combining the individual fitting results, and our algorithm offers an interpretable solution.
@@ 3.1 W51-Mopra
We utilize the published data from Watanabe et al. (2017), who conducted a line survey towards the W51 region using the Mopra 22 m telescope. The survey covers the frequency ranges from 85.2 - 101.1 GHz and 107.0 - 114.9 GHz. The spectral data originate from the hot cores e1/e2, located at galactic coordinate ($l$, $b$) = (49.4898, -0.3874). Watanabe et al. (2017) identified 234 emission lines. By matching the spectral line databases, namely CDMS and JPL, they identified 31 molecules and 18 isotopologues. No parameter estimation is provided for the spectrum of the hot cores e1/e2.
@@ 3.2 OrionKL-GBT
This study relies on the data published by Frayer et al. (2015), who conducted a line survey towards Orion KL. The observations were carried out using the Green Bank Telescope (GBT) and were centered on the celestial coordinates R.A. (J2000) 05:35:14, decl. (J2000) 05:22:27.5. The frequency range covered by the survey spans from 67.0 GHz to 93.6 GHz. Frayer et al. (2015) only identified lines with intensities brighter than 1 K. By matching the Splatalogue database, a total of 140 emission lines were identified, with no parameter estimation provided.
@@ 3.3 OrionKL-Tianma
We incorporate the data from Liu et al. (2022), who carried out a line survey targeting the Orion KL region using the Tianma 65 m radio telescope. The spectral data are centered on the celestial coordinates R.A. (J2000) 05:35:14.55, decl. (J2000) -05:22:31.0. The frequency coverage spans from 34.5 GHz to 50.0 GHz. A total of 597 emission lines were identified, including 177 classified as radio recombination lines. Liu et al. (2022) carried out a spectral line fitting analysis for the molecular lines using a one-dimensional LTE model. Their spectral line model employs equations similar to Equations 1 to 3. In their analysis, they adopted fixed values for the source size and excitation temperature, while fitting the column density, velocity width, and velocity offset.
@@ 3.4 SgrB2M-IRAM and SgrB2N-IRAM
The most complex test data utilized in this study originate from Belloche et al. (2013), who presented a line identification analysis for both Sgr B2(M) and B2(N) within the frequency range of 80.0 GHz to 116.0 GHz. The observations were conducted using the IRAM 30 m telescope, targeting the positions at R.A. (J2000) 17:47:20s.4, decl. (J2000) -28:23:07.0 for Sgr B2(M), and R.A. (J2000) 17:47:20.0 and decl. (J2000) -28:22:19.0 for and B2(N), respectively. These sources are line rich. Belloche et al. (2013) identified 945 and 3675 emission lines in Sgr B2(M) and B2(N) respectively. They fit the observed spectra using the xclass spectral model to estimate the source size, column density, excitation temperature, velocity width and velocity offset. The parameters were adjusted through visual inspection.
@@
Name Band Line densitya Recallb ($\chi^{2}$) Recallb (PM) Prec.c ($\chi^{2}$) Prec.c (PM)
(GHz) (GHz-1)
W51-Mopra 85 - 101, 107 - 115 13.8 61% (93/152) 84% (127/152) 73% 90%
OrionKL-GBT 67 - 94 20.3 82% (86/105) 93% (98/105) 86% 89%
OrionKL-Tianma 35 - 50 30.5 57% (136/237) 83% (196/237) 62% 89%
SgrB2M-IRAM 80 - 116 38.0 39% (181/463) 78% (363/463) 56% 92%
SgrB2N-IRAM 80 - 116 59.2 31% (329/1055) 74% (784/1055) 51% 78%
Figure 4: Comparison of the best-fitting spectra using different loss functions. The black solid line illustrates the observed spectrum of SgrB2M-IRAM (see Section 3.4), and blue dashed lines represent the spectral lines of C2H3CN identified by Belloche et al. (2013). The blue and red lines demonstrate the best-fitting results achieved with the $\chi^{2}$ and peak matching loss functions respectively. The peak matching loss function is defined in Section 2.4. We compute and display both the $\chi^{2}$ and peak matching loss values corresponding to each best-fitting spectrum. These values imply that using the $\chi^{2}$ loss is inappropriate for spectral fitting of a single molecule.